OpenStudy (anonymous):

An iron ball 4 inches in diameter weighs 9 pounds. Find the Weight of an iron shell 2 inches thick whose external diameter is 20 inches.

OpenStudy (anonymous):

First you need to find how much one cubic centimeter of iron weighs

OpenStudy (anonymous):

So you would set it up like\[\frac{ 4\pi(2)^2 }{ 9}\]

OpenStudy (anonymous):

to find lb/cubic inches

OpenStudy (anonymous):

and then the iron shell is a hemi sphere

OpenStudy (anonymous):

can you explain it step by step? I don't seem to get it. :/

OpenStudy (anonymous):

\[\frac{ area~of~sphere }{ weight~of~sphere }= Pounds~per~cubic~inch\]

OpenStudy (anonymous):

\[2\pi(2)^2=Cubic~inches~inside~of~hemisphere\]

OpenStudy (anonymous):

\[Pounds~per~cubic~inch*area~of~hemisphere=weight~of~hemisphere\]

OpenStudy (anonymous):

is it area or volume of the sphere? -.-